A mixture of gas consists of 10 g N_2, 10 g O_2, a
Answer #1

no of moles of N2   nN2     = W/G.M.Wt   = 10/28   = 0.357 moles

no of moles of O2   nO2     = W/G.M.Wt    = 10/32   = 0.3125 moles

no of moles of C2H2   nC2H2     = W/G.M.Wt    = 10/26 = 0.384 moles

mole fraction of XN2      = nN2 /nN2 + nO2 + nC2H2

                                      = 0.375/0.375+0.3125+0.384

                                       = 0.375/1.0715 = 0.349

mole fraction of XO2      = nO2 /nN2 + nO2 + nC2H2

                                      = 0.3125/0.375+0.3125+0.384

                                       = 0.3125/1.0715    = 0.292

mole fraction of XC2H2      = nC2H2 /nN2 + nO2 + nC2H2

                                      = 0.384/0.375+0.3125+0.384

                                       = 0.384/1.0715      = 0.358

D. >>> answer